OpenStudy (idku):

Please help me with 2 brief questions.

OpenStudy (idku):

A cannonball is shot (from ground level) with an initial horizontal velocity of 33 m/s and an initial vertical velocity of 25 m/s. ------------------------------------------------------------ \(\rm \large I~found~so~far:\) [1] initial speed, \(\rm u=41.4(m/s)\) [2] angle of cannonball with respect to ground, \(\rm \theta=37.15º\) [3] maximum height the cannonball goes above the ground \(\rm x=31.8(m)\) [4] horizontal displacement is \(\rm 167.6(m)\) [5] the entire time of the ball motion \(\rm 5.1(s)\) ------------------------------------------------- \(\rm \large I~need:\) [6] speed of the cannonball at, \(\rm t=3.7(s)\) [7] height (Vertical Displacement) at, \(\rm t=3.7(s)\)

OpenStudy (anonymous):

6.Vi = √[Vx²+Vy²] = 44.2 m/s 7.3) Hmax = Vi²sin²Θ/(2g) = 62.49 m

OpenStudy (idku):

Are you sure? Becsause that doesn't seem to be right.

OpenStudy (anonymous):

Don't know,I never liked physics. Maybe he can help @Hero @dan815 @iambatman

OpenStudy (anonymous):

@triciaal

OpenStudy (idku):

I found the vertical displacement x=25.35m applying: \(x = ut + \dfrac{1}{2}at^2\) I have put: t=3.7s a=-9.81m/s\(^2\) u=25m/s

OpenStudy (anonymous):

Nice.

OpenStudy (idku):

I need velocity at t=3.7 the rest i found. tnx for coming btw

OpenStudy (anonymous):

Your welcome!

OpenStudy (dan815):

A cannonball is shot (from ground level) with an initial horizontal velocity of 33 m/s and an initial vertical velocity of 25 m/s. [6] speed of the cannonball at, t=3.7(s) [7] height (Vertical Displacement) at, t=3.7(s)

OpenStudy (idku):

[7]=25.35m found that:)

OpenStudy (dan815):